OpenStudy (madgirlwithabluebox):

can someone walk me through this? what is the sum of the finite arithmetic series? (-15)+0+15+30...+195

OpenStudy (solomonzelman):

okay, what is the difference in the sequence ?

OpenStudy (madgirlwithabluebox):

+15

OpenStudy (solomonzelman):

yes.

OpenStudy (solomonzelman):

So, you add 15 every time

OpenStudy (solomonzelman):

And you start from -15

OpenStudy (solomonzelman):

Can you think of a line with a slope of 15 (common difference 15) and a y-intercept (-15)(the first term) ?

OpenStudy (solomonzelman):

what will the equation of this line be ?

OpenStudy (madgirlwithabluebox):

uh what?

OpenStudy (solomonzelman):

y-intercept (or the first term) = -15 difference (or the slope) = 15 what is the equation of the line ?

OpenStudy (solomonzelman):

(I am just trying to get you to think of a pattern that models the terms, and then we will do the sigma notation)

OpenStudy (madgirlwithabluebox):

y=15x-15?

OpenStudy (solomonzelman):

yes, exactly

OpenStudy (solomonzelman):

now, we need to see whichTH term is 195

OpenStudy (solomonzelman):

\(\large\color{blue}{ \displaystyle a_n=a_1+d(n-1)}\) \(\large\color{blue}{ \displaystyle 195=(-15)+15(n-1)}\) see how I am setting it up? Can you solve for "n" ? (this n-solution is going to be the number of term that 195 has in the sequence)

OpenStudy (madgirlwithabluebox):

196=n?

OpenStudy (solomonzelman):

how so ?

OpenStudy (madgirlwithabluebox):

oh wait

OpenStudy (madgirlwithabluebox):

i see what i did

OpenStudy (madgirlwithabluebox):

11=n?

OpenStudy (solomonzelman):

the 15 (i plugged for a1) is negative and you are adding n to both sides at the end

OpenStudy (solomonzelman):

I mean adding 1 to both sides at the end

OpenStudy (solomonzelman):

creative way \(\large\color{blue}{ \displaystyle 195=(-15)+15(n-1)}\) \(\large\color{blue}{ \displaystyle 195=15(n-1-1)}\) \(\large\color{blue}{ \displaystyle 195=15(n-2)}\) \(\large\color{blue}{ \displaystyle 13=n-2}\)

OpenStudy (solomonzelman):

or, same way you can add 15 to both sides, and the divide both sides by 15, you would be getting \(\large\color{blue}{ \displaystyle 14=n-1}\)

OpenStudy (madgirlwithabluebox):

well let me show you how i did it so i can see what i did wrong

OpenStudy (solomonzelman):

sure

OpenStudy (madgirlwithabluebox):

\[195=(-15)+15(n-1)\] \[195=(-15)+15n-15\] \[195=-30+15n\] (i realized what i did wrong)..

OpenStudy (madgirlwithabluebox):

i subtracted 30 instead of added it

OpenStudy (madgirlwithabluebox):

225/15=15n/15

OpenStudy (madgirlwithabluebox):

15=n

OpenStudy (solomonzelman):

yes. and sorry an accident happened

OpenStudy (solomonzelman):

My young 12 year old brother plugged in a bare wire, into the plugin, and somewhere in a room I am sitting, and my comp shut down completely. Fortunately, no one suffered of electric current...:) Anyways

OpenStudy (solomonzelman):

so you have 15 terms from n=1 to n=15 OR from n=0 to n=14

OpenStudy (solomonzelman):

\(\large\color{blueviolet}{ \displaystyle \sum_{ n=0 }^{ 14 } ~(15n-15)}\)

OpenStudy (madgirlwithabluebox):

ohkay how do i begin with that

OpenStudy (solomonzelman):

well, what I wrote is one possible answer (and I believe the most simple one

OpenStudy (madgirlwithabluebox):

ell the answers its giving me are 1355 1350 1335 1000

OpenStudy (madgirlwithabluebox):

i believe its 1350 but im not completely sure

OpenStudy (solomonzelman):

I http://www.wolframalpha.com/input/?i=sum+from+n%3D0+to+n%3D14+%2815n-15%29 correct!